# Query Performance Prediction using Relevance Judgments Generated by Large Language Models

### v2

@@ 2.2. Zero/few-shot prompting and parameter-efficient fine-tuning (PEFT) for LLMs
Automatically generating relevance judgments is a long-standing goal in IR that has been studied for multiple decades (Makary et al., 2017, 2016; Ravana et al., 2015; Nuray and Can, 2006, 2003; Soboroff et al., 2001). Recent studies have demonstrated promising results of using LLMs for the automatic generation of relevance judgments (Thomas et al., 2023; Faggioli et al., 2023a). In this paper we focus on studies into generating relevance judgments with discrete classes (e.g., “Relevant” or “Irrelevant”), instead of generating continuous relevance labels in real numbers (Yan et al., 2024). We discuss related studies into LLM-based automatic generation of relevance judgments from two dimensions: (i) how LLMs are used to generate relevance judgments, and (ii) their applications.
@@
Recent studies have explored prompting commercial LLMs (e.g., GPT-3.5 and GPT 4) or open-source LLMs in zero- or few-shot manners. Specifically, Faggioli et al. (2023a) use zero- and few-shot prompting to instruct GPT-3.5 to predict the relevance of an item to a query. Thomas et al. (2023) instruct GPT-4 by zero-shot prompting, and add to the prompt a detailed query description and consider chain-of-thought (Wei et al., 2022). Similarly, Rahmani et al. (2024) follow Thomas et al. (2023)’s prompt to instruct GPT-4. Ma et al. (2024) instruct GPT-3.5 to generate relevance judgments for a domain-specific scenario, i.e., legal case retrieval (Ma et al., 2021); they use prompts specifically designed for this scenario. More recently, Upadhyay et al. (2024b) prompt GPT-4o in a zero-shot manner. Besides using commercial LLMs, only few studies (Zhang et al., 2024; Salemi and Zamani, 2024; Upadhyay et al., 2024a; Khramtsova et al., 2024) explore prompting open-source LLMs to generate relevance judgments. E.g., Khramtsova et al. (2024), Upadhyay et al. (2024a) and Salemi and Zamani (2024) prompt Flan-T5 (Chung et al., 2022), Vicuña-7B (Zheng et al., 2024) and Mistral (Jiang et al., 2023), respectively, in either zero-shot or few-shot manners. MacAvaney and Soldaini (2023) focus on a special scenario where a relevant item for a given query is already known and use Flan-T5 (Chung et al., 2022) to estimate the relevance of another item to the query given the known relevant item.
@@
Recent studies have explored using LLM-generated relevance judgments to benefit (i) search system evaluation (Upadhyay et al., 2024a; Rahmani et al., 2024; Thomas et al., 2023; MacAvaney and Soldaini, 2023; Faggioli et al., 2023a), (ii) ranker selection (Khramtsova et al., 2024), (iii) item selection and retrieval quality evaluation in retrieval-augmented generation (RAG(Zhang et al., 2024; Salemi and Zamani, 2024) and (iv) retriever fine-tuning (Ma et al., 2024). Concerning (i), recent studies (Rahmani et al., 2024; Upadhyay et al., 2024a; Abbasiantaeb et al., 2024; Faggioli et al., 2023a; Thomas et al., 2023; MacAvaney and Soldaini, 2023) explore evaluating search systems either entirely using LLM-generated relevance judgments or partially using LLM-generated relevance judgments (a.k.a. filling holes). They have demonstrated a high correlation between search system rankings based on LLM- and human-labelled relevance judgments. As to (ii), given a pool of dense retrievers, Khramtsova et al. (2024) select a suitable one for a target corpus by estimating their performance using LLM-generated queries and relevance judgments specific to the target corpus. For (iii), for item selection, Zhang et al. (2024) prompt LLMs to generate relevance judgments for retrieved candidate items in RAG; the items that are predicted as “relevant” are used for text generation. Zhang et al. (2024) observe that items selected via relevance prediction resulted in sub-optimal text generation quality. For retrieval quality evaluation, Salemi and Zamani (2024) generate relevance judgments for retrieved candidate items and aggregate those judgments into a score. However, Salemi and Zamani (2024) found that the aggregated score based on the LLM-generated relevance judgments achieves a low correlation with the text generation quality of RAG. Concerning (iv), Ma et al. (2024) fine-tune a legal case retriever on a training set augmented with LLM-generated relevance judgments. They show that fine-tuning a legal case retriever using the generated relevance judgments results in enhanced performance.
@@
Our main contribution in this paper is the introduction of QPP-GenRE, a novel QPP framework, which, in theory, can be integrated with various relevance prediction approaches. To demonstrate the compatibility of QPP-GenRE with various relevance prediction approaches, we adapt a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a), into a relevance judgment generator by applying a threshold for its re-ranking scores; we then integrate QPP-GenRE with this adapted RankLLaMA. Note that adapting other types of LLM-based re-rankers, e.g., pairwise and listwise LLM-based re-rankers into relevance judgment generators is challenging because they do not explicitly produce explicit relevance scores. Exploring their use as relevance judgment generators is beyond the scope of this paper.
@@ 3. Task definition
(1) $p=\psi(q,L,C)\in\mathbb{R}~{},$
where $p$ indicates the predicted retrieval quality of the ranker $M$ in response to the query $q$; typically, $p$ is expected to be correlated with an IR evaluation measure, such as reciprocal rank (RR).
@@ 4.1. Overview of QPP-GenRE
We propose QPP-GenRE, which consists of two steps: (i) generating relevance judgments using LLMs, and (ii) predicting IR evaluation measures. In (i), we employ an LLM to generate relevance judgments for the top-$n$ retrieved items in the ranked list for a given query; to improve LLM’s effectiveness in generating relevance judgments, we fine-tune an LLM with PEFT using human-labeled relevance judgments. In (ii), we regard the generated relevance judgments as pseudo labels to calculate different IR evaluation measures.
@@ 4.2.1. Inference
(2) $\hat{r}_{i}=\mathrm{LLM}(q,d_{i})~{},\vspace*{-1.5mm}$
where $\hat{r}_{i}$ is a predicted relevance value for the item $d_{i}$ at rank $i$. $\hat{r}_{i}\in\{1,0\}$ , where “1” indicates relevant and “0” irrelevant. We leave the prediction of multi-graded labels as future work. After automatically judging the top-$n$ items in the ranked list $L$, we get a list of generated relevant judgments ^$\hat{\mathcal{R}}_{L_{1:n}}=[\hat{r}_{1},\dots,\hat{r}_{i},\dots,\hat{r}_{n}]$, where $\hat{r}_{i}$ is the predicted relevance value for $d_{i}$ in $L$. We design a prompt to instruct an LLM on the task of automatic generation of relevance judgments, as illustrated in Figure 1.
@@
To further improve an LLM’s effectiveness in generating relevance judgments, we use human-labeled relevance judgments to train an LLM with an effective PEFT method, QLoRA (Dettmers et al., 2023). Specifically, we first quantize an LLM model to 4-bit, add learnable low-rank adapters to each network layer of the LLM, and then optimize low-rank adapters. Formally, given the query $q$ and an item $d_{i}$ in the ranked list $L$, we optimize the LLM to generate the human-labeled relevance value $r_{i}$ for the item $d_{i}$. See Section 5 for more details.
@@ 4.3.1. Predicting precision-oriented measures
We compute a precision-oriented measure based on LLM generated relevant judgments ^$\hat{\mathcal{R}}_{L_{1:n}}$ for the top-$n$ items in the ranked list $L$. The following is an example to compute RR@k:
(3) $RR@k=1/\min_{i}\{\hat{r}_{i}>0\}~{},$
where $0<i\leqslant k$. $RR@k$ would be equal to 0 if there is no top-$k$ item that is predicted as relevant to the query $q$; in this case, $n=k$.
@@ 4.3.2. An approximation strategy to predict measures considering recall
As the computation of a measure considering recall requires the information of all relevant items in the corpus $C$ for a given query $q$, we need to automatically assess every item in corpus $C$, which is infeasible due to the high computational cost. To address this issue, we devise an approximation strategy for predicting an IR measure considering recall, which only judges the top-$n$ items in the ranked list $L$ and uses the items predicted as relevant to approximate all relevant items in the corpus, to avoid the cost of judging the entire corpus. Fröbe et al. (2023); Moffat (2017); Lu et al. (2016) define normalized discounted cumulative gain (nDCG(Järvelin and Kekäläinen, 2002) at a cutoff $k$ as a recall-oriented IR evaluation metric because it is normalized by a recall-oriented “best possible” ranking. [^1] In this paper, we employ nDCG@10 and believe that nDCG@10 is a metric considering recall: Figure 3 illustrates that to reach saturation in predicting nDCG@10 values for ANCE and BM25, judgments up to the top 100 and 200 retrieved items are needed, respectively. If it were a precision-based metric, saturation could be achieved by judging around 10 items. nDCG@10 is also the most primary official IR evaluation metric in TREC-DL 19–22 (Craswell et al., 2022, 2021, 2020, 2019). Thus, here we show an example of predicting nDCG@$k$ (Järvelin and Kekäläinen, 2002), formally:
(4) $\begin{split}nDCG@k={DCG@k}/{IDCG@k}~{},\end{split}$
@@
(5) $\begin{split}DCG@k=\hat{r}_{1}+\sum_{i=2}^{k}\hat{r}_{i}/\log_{2}i~{}.\end{split}$
@@
(6) $\begin{split}IDCG@k=\hat{ir}_{1}+\sum_{i=2}^{k}\hat{ir}_{i}/\log_{2}i~{}.\end{split}\vspace*{-1.5mm}$
@@ 2. RQ2
    To what extent does QPP-GenRE improve QPP quality for lexical and neural rankers in terms of nDCG@10, an IR metric considering recall, compared to state-of-the-art baselines?
@@ 3. RQ3
    How does judging depth in a ranked list affect the prediction of nDCG@10, an IR metric that considers recall?
@@ 5.3. Retrieval approaches
We consider BM25 (Robertson et al., 2009) as a lexical ranker, and ANCE (Xiong et al., 2021) as a neural-based dense retriever. To increase the comparability and reproducibility of our paper, we get the retrieval results of both rankers using the publicly available resource from Pyserini (Lin et al., 2021). We get BM25’s retrieval result with top-1000 retrieved items per query on the TREC-DL 19–22 datasets using the default parameters ($k1=0.9$, $b=0.4$). BM25’s actual nDCG@10 values are 0.506, 0.480, 0.446 and 0.269 on TREC-DL 19, 20, 21 and 22, respectively. We get ANCE’s retrieval result with top-1000 retrieved items per query on TREC-DL 19–20, using the publicly available dense vector index of ANCE on MS MARCO V1. ANCE’s actual nDCG@10 values are 0.645 and 0.646 on TREC-DL 19 and 20, respectively. We rely on the publicly available dense vector index of ANCE; at the time of writing, there is no dense vector index of ANCE publicly available on MS MARCO V2 for TREC-DL 21 and 22.333Building the dense vector index on MS MARCO V2 with over 138 million passages is resource-intensive and out of the scope of our work.
@@ 5.5. QPP evaluation and target IR evaluation measures
We follow established best practices (Datta et al., 2022c; Zamani et al., 2018; Carmel and Yom-Tov, 2010; Hauff et al., 2008; Cronen-Townsend et al., 2002) to evaluate QPP by measuring linear correlation by Pearson’s $\rho$ as well as ranked-based correlation through Kendall’s $\tau$ correlation coefficients between the actual and predicted performance of a query set. [^4] We also consider scaled Mean Absolute Ranking Error (sMARE) (Faggioli et al., 2022, 2021b) and draw the same conclusion as Pearson’s $\rho$ and Kendall’s $\tau$; we include the results in our repository at https://github.com/ChuanMeng/QPP-GenRE. As for target IR metrics, we consider the two primary official IR metrics used in TREC DL 19–22 (Craswell et al., 2022, 2021, 2020, 2019), RR@10 (precision-oriented) and nDCG@10 (considering recall); recent QPP studies (Arabzadeh et al., 2023; Faggioli et al., 2023e; Khodabakhsh and Bagheri, 2023) consider either or both of these metrics as their target metrics. Following (Datta et al., 2022c), we use relevance scale $\geq$ 2 as positive to compute actual binary IR measures (e.g., RR). When calculating correlation for nDCG@10, the actual values of nDCG@10 are calculated by human-labeled and multi-graded relevance judgments, while the nDCG@10 values predicted by QPP-GenRE are based on its generated binary judgments.

### v3

@@ 2.3. LLMs for generating relevance judgments
Automatically generating relevance judgments is a long-standing goal in IR that has been studied for multiple decades (Makary et al., 2017, 2016; Ravana et al., 2015; Nuray and Can, 2006, 2003; Soboroff et al., 2001). Recent studies have demonstrated promising results of using LLMs for the automatic generation of relevance judgments (Thomas et al., 2024; Faggioli et al., 2023a). In this paper we focus on studies into generating relevance judgments with discrete classes (e.g., “Relevant” or “Irrelevant”), instead of generating continuous relevance labels in real numbers (Yan et al., 2024). We discuss related studies into LLM-based automatic generation of relevance judgments from two dimensions: (i) how LLMs are used to generate relevance judgments, and (ii) their applications.
@@
Recent studies have explored prompting commercial LLMs (e.g., GPT-3.5 and GPT 4) or open-source LLMs in zero- or few-shot manners. Specifically, Faggioli et al. (2023a) use zero- and few-shot prompting to instruct GPT-3.5 to predict the relevance of an item to a query. Thomas et al. (2024) instruct GPT-4 by zero-shot prompting, and add to the prompt a detailed query description and consider chain-of-thought (Wei et al., 2022). Ma et al. (2024) instruct GPT-3.5 to generate relevance judgments for a domain-specific scenario, i.e., legal case retrieval (Ma et al., 2021); they use prompts specifically designed for this scenario. More recently, Upadhyay et al. (2024c) prompt GPT-4o in a zero-shot manner. Besides using commercial LLMs, only few studies (Zhang et al., 2024; Salemi and Zamani, 2024; Upadhyay et al., 2024a; Khramtsova et al., 2024) explore prompting open-source LLMs to generate relevance judgments. E.g., Khramtsova et al. (2024), Upadhyay et al. (2024a) and Salemi and Zamani (2024) prompt Flan-T5 (Chung et al., 2022), Vicuña-7B (Zheng et al., 2024) and Mistral (Jiang et al., 2023), respectively, in either zero-shot or few-shot manners. MacAvaney and Soldaini (2023) focus on a special scenario where a relevant item for a given query is already known and use Flan-T5 (Chung et al., 2022) to estimate the relevance of another item to the query given the known relevant item.
@@
Recent studies have explored using LLM-generated relevance judgments to benefit (i) search system evaluation (Upadhyay et al., 2024a; Thomas et al., 2024; MacAvaney and Soldaini, 2023; Faggioli et al., 2023a), (ii) ranker selection (Khramtsova et al., 2024), (iii) item selection and retrieval quality evaluation in retrieval-augmented generation (RAG(Zhang et al., 2024; Salemi and Zamani, 2024) and (iv) retriever fine-tuning (Ma et al., 2024). Concerning (i), recent studies (Upadhyay et al., 2024a; Abbasiantaeb et al., 2024; Faggioli et al., 2023a; Thomas et al., 2024; MacAvaney and Soldaini, 2023) explore evaluating search systems either entirely using LLM-generated relevance judgments or partially using LLM-generated relevance judgments (a.k.a. filling holes). They have demonstrated a high correlation between search system rankings based on LLM- and human-labelled relevance judgments. As to (ii), given a pool of dense retrievers, Khramtsova et al. (2024) select a suitable one for a target corpus by estimating their performance using LLM-generated queries and relevance judgments specific to the target corpus. For (iii), for item selection, Zhang et al. (2024) prompt LLMs to generate relevance judgments for retrieved candidate items in RAG; the items that are predicted as “relevant” are used for text generation. Zhang et al. (2024) observe that items selected via relevance prediction resulted in sub-optimal text generation quality. For retrieval quality evaluation, Salemi and Zamani (2024) generate relevance judgments for retrieved candidate items and aggregate those judgments into a score. However, Salemi and Zamani (2024) found that the aggregated score based on the LLM-generated relevance judgments achieves a low correlation with the text generation quality of RAG. Concerning (iv), Ma et al. (2024) fine-tune a legal case retriever on a training set augmented with LLM-generated relevance judgments. They show that fine-tuning a legal case retriever using the generated relevance judgments results in enhanced performance.
@@ 2.4. LLMs for re-ranking
Our main contribution in this paper is the introduction of QPP-GenRE, a novel QPP framework, which, in theory, can be integrated with various relevance prediction approaches. To demonstrate the compatibility of QPP-GenRE with various relevance prediction approaches, we adapt a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a), into a relevance judgment generator by applying a threshold for its re-ranking scores; we then integrate QPP-GenRE with this adapted RankLLaMA. It is important to note that exploring the use of other types of LLM-based re-rankers (e.g., pairwise and listwise) as relevance judgment generators falls outside the scope of this paper.
@@ 3. Task definition
(1) $p=\psi(q,L,C)\in\mathbb{R},$
where $p$ indicates the predicted retrieval quality of the ranker $M$ in response to the query $q$; typically, $p$ is expected to be correlated with an IR evaluation measure, such as RR.
(a) Predicting a precision-based metric. In this case, we predicts the top-$n$ items in the given ranked list, $n$ is equal to $k$, which is the cut-off point for a precision-oriented metric, such as reciprocal rank (RR). E.g., the predicted $RR@k$ for the ranked list in this figure is 0.5 as the second item is predicted as relevant.
(b) Predicting a metric considering recall using our proposed approximation strategy. It only judges the top-$n$ ($n$ $\ll$ the total number of items in the corpus)
items in the given ranked list and uses the items predicted as relevant to approximate all relevant items in the corpus.
Figure 1. The framework of QPP-GenRE.
@@ 4.1. Overview of QPP-GenRE
We propose QPP-GenRE, which consists of two steps: (i) generating relevance judgments using LLMs, and (ii) predicting IR evaluation measures. See Fig. 1. In (i), we employ an LLM to generate relevance judgments for the top-$n$ retrieved items in the ranked list for a given query; to improve LLMs’s effectiveness in generating relevance judgments, we fine-tune an LLM with PEFT using human-labeled relevance judgments. In (ii), we regard the generated relevance judgments as pseudo labels to calculate different IR evaluation measures.
@@ 4.2.1. Inference
(2) $\hat{r}_{i}=\mathrm{LLM}(\mathrm{prompt}(q,d_{i})),$
where $\mathrm{prompt}(\cdot,\cdot)$ is a prompt to instruct an LLM on the task of automatic generation of relevance judgments, as illustrated in Figure 2. We follow the design proposed by Faggioli et al. (2023a) to create a prompt that explicitly instructs the LLM to output either “Relevant” or “Irrelevant”. In our preliminary experiments, we also tested the prompt from Sun et al. (2023b), which asks the LLM the question, “Does the passage answer the query?” and expects a response of either “Yes” or “No.” However, we found that this alternative prompt produced inferior results compared to our chosen design. $\hat{r}_{i}$ is a predicted relevance value for the item $d_{i}$ at rank $i$. $\hat{r}_{i}\in\{1,0\}$ , where “1” indicates relevant and “0” irrelevant. We leave the prediction of multi-graded labels as future work. After automatically judging the top-$n$ items in the ranked list $L$, we get a list of generated relevant judgments ^$\hat{\mathcal{R}}_{L_{1:n}}=[\hat{r}_{1},\dots,\hat{r}_{i},\dots,\hat{r}_{n}]$, where $\hat{r}_{i}$ is the predicted relevance value for $d_{i}$ in $L$.
@@
To further improve an LLM’s effectiveness in generating relevance judgments, we use human-labeled relevance judgments to train an LLM with an effective PEFT method, QLoRA (Dettmers et al., 2023). Specifically, we first quantize an LLM model to 4-bit, add learnable low-rank adapters to each network layer of the LLM, and then optimize low-rank adapters. Formally, given the query $q$ and an item $d_{i}$ in the ranked list $L$, we optimize the LLM to generate the human-labeled relevance value $r_{i}$ for the item $d_{i}$:
(3) $\begin{split}\mathcal{L}(\theta_{LoRA})=-\frac{1}{M}\sum_{i=1}^{M}\log P(r_{i}\mid\mathrm{prompt}(q,d_{i})),\end{split}$
where $\theta_{LoRA}$ stands for learnable low-rank adapters added to the LLM; $M$ is the number of training examples. See Section 5.6 for more details.
@@ 4.3.1. Predicting precision-oriented measures
We compute a precision-oriented measure based on LLM generated relevant judgments ^$\hat{\mathcal{R}}_{L_{1:n}}$ for the top-$n$ items in the ranked list $L$, as shown in Figure 1(a). Note that in this case, $n=k$. The following is an example to compute RR@k:
(4) $RR@k=1/\min_{i}\{\hat{r}_{i}>0\},$
where $0<i\leqslant k$. For instance, as illustrated in Figure 1(a), the first item in the ranked list is predicted as irrelevant, while the second item is predicted as relevant. In this case, the predicted $RR@k$ value would be 0.5. $RR@k$ would be equal to 0 if there is no top-$k$ item that is predicted as relevant to the query $q$.
@@ 4.3.2. An approximation strategy to predict measures considering recall
As the computation of a measure considering recall requires the information of all relevant items in the corpus $C$ for a given query $q$, we need to automatically assess every item in corpus $C$, which is infeasible due to the high computational cost. To address this issue, we devise an approximation strategy for predicting an IR measure considering recall, which only judges the top-$n$ ($n$ $\ll$ the total number of items in the corpus) items in the ranked list $L$ and uses the items predicted as relevant to approximate all relevant items in the corpus, to avoid the cost of judging the entire corpus. Fröbe et al. (2023); Moffat (2017); Lu et al. (2016) define normalized discounted cumulative gain (nDCG(Järvelin and Kekäläinen, 2002) at a cutoff $k$ as a recall-oriented IR evaluation metric because it is normalized by a recall-oriented “best possible” ranking. [^1] In this paper, we employ nDCG@10 and believe that nDCG@10 is a metric considering recall: Figure 4 illustrates that to reach saturation in predicting nDCG@10 values for ANCE and BM25, judgments up to the top 100 and 200 retrieved items are needed, respectively. If it were a precision-based metric, saturation could be achieved by judging around 10 items. nDCG@10 is also the most primary official IR evaluation metric in TREC-DL 19–22 (Craswell et al., 2022, 2021, 2020, 2019). Thus, here we show an example of predicting nDCG@$k$ (Järvelin and Kekäläinen, 2002), formally:
(5) $\begin{split}nDCG@k={DCG@k}/{IDCG@k},\end{split}$
@@
(6) $\begin{split}DCG@k=\hat{r}_{1}+\sum_{i=2}^{k}\hat{r}_{i}/\log_{2}i.\end{split}$
@@
(7) $\begin{split}IDCG@k=\hat{ir}_{1}+\sum_{i=2}^{k}\hat{ir}_{i}/\log_{2}i.\end{split}$
@@ 2. RQ2
    To what extent does QPP-GenRE improve QPP quality for lexical and neural rankers in terms of nDCG@10, an IR metric that not only considers precision but also takes recall into account, compared to state-of-the-art baselines?
@@ 3. RQ3
    How does judging depth in a ranked list affect the prediction of nDCG@10, an IR metric that considers both precision and recall? In other words, how does varying the number of top-ranked documents submitted for relevance judgments impact QPP quality?
@@ 5.3. Retrieval approaches
We consider BM25 (Robertson et al., 1995) as a lexical ranker; we also consider ANCE (Xiong et al., 2021) and TAS-B (Hofstätter et al., 2021) as neural-based dense retrievers. To increase the comparability and reproducibility of our paper, we get the retrieval results of both rankers using the publicly available resource from Pyserini (Lin et al., 2021). We get BM25’s retrieval result with top-1000 retrieved items per query on the TREC-DL 19–22 datasets using the default parameters ($k1=0.9$, $b=0.4$). BM25’s actual nDCG@10 values are 0.506, 0.480, 0.446 and 0.269 on TREC-DL 19, 20, 21 and 22, respectively. We get the retrieval results of ANCE and TAS-B with top-1000 retrieved items per query on TREC-DL 19–20, using the publicly available dense vector index of ANCE on MS MARCO V1. ANCE’s actual nDCG@10 values are 0.645 and 0.646 on TREC-DL 19 and 20, respectively; TAS-B’s actual nDCG@10 values are 0.721 and 0.685 on TREC-DL 19 and 20, respectively. We rely on the publicly available dense vector index of ANCE/TAS-B; at the time of writing, there is no dense vector index of ANCE/TAS-B publicly available on MS MARCO V2 for TREC-DL 21 and 22.333Building the dense vector index on MS MARCO V2 with over 138 million passages is resource-intensive and beyond the scope of our work.
@@ 5.5. QPP evaluation and target IR evaluation measures
We follow established best practices (Datta et al., 2022c; Zamani et al., 2018; Carmel and Yom-Tov, 2010; Hauff et al., 2008; Cronen-Townsend et al., 2002) to evaluate QPP by measuring linear correlation by Pearson’s $\rho$ as well as ranked-based correlation through Kendall’s $\tau$ correlation coefficients between the actual and predicted performance of a query set. As for target IR metrics, we consider the two primary official IR metrics used in TREC DL 19–22 (Craswell et al., 2022, 2021, 2020, 2019), RR@10 (precision-oriented) and nDCG@10 (considering recall); recent QPP studies (Arabzadeh et al., 2023; Faggioli et al., 2023e; Khodabakhsh and Bagheri, 2023) consider either or both of these metrics as their target metrics. Following (Datta et al., 2022c), we use relevance scale $\geq$ 2 as positive to compute actual binary IR measures (e.g., RR). When calculating correlation for nDCG@10, the actual values of nDCG@10 are calculated by human-labeled and multi-graded relevance judgments, while the nDCG@10 values predicted by QPP-GenRE are based on its generated binary judgments.